# Large Language Models for Generative Information Extraction: A Survey

### v2

@@ Abstract
Information extraction (IE) aims to extract structural knowledge from plain natural language texts. Recently, generative Large Language Models (LLMs) have demonstrated remarkable capabilities in text understanding and generation. As a result, numerous works have been proposed to integrate LLMs for IE tasks based on a generative paradigm. To conduct a comprehensive systematic review and exploration of LLM efforts for IE tasks, in this study, we survey the most recent advancements in this field. We first present an extensive overview by categorizing these works in terms of various IE subtasks and techniques, and then we empirically analyze the most advanced methods and discover the emerging trend of IE tasks with LLMs. Based on a thorough review conducted, we identify several insights in technique and promising research directions that deserve further exploration in future studies. We maintain a public repository and consistently update related resources  [^1] https://github.com/quqxui/Awesome-LLM4IE-Papers.
@@
†Equal Contribution.,
@@ 1 Introduction
Information Extraction (IE) is a crucial domain in natural language processing (NLP) that converts plain text into structured knowledge (e.g., entities, relations, and events), and serves as a foundational requirement for a wide range of downstream tasks, such as knowledge graph construction [245], knowledge reasoning [51] and question answering [180]. Typical IE tasks consist of Named Entity Recognition (NER), Relation Extraction (RE) and Event Extraction (EE) [133, 204, 64, 175]. On the other hand, the emergence of large language models (LLMs) (e.g., GPT-4 [1]) has greatly promoted the development of NLP, due to their extraordinary capabilities in text understanding and generation. Therefore, there has been a recent surge of interest in generative IE methods [166, 175] that adopt LLMs to generate structural information rather than extracting structural information from plain text. These methods have been proven to be more practical in real-world scenarios compared to discriminated methods [24, 128], as they efficiently handle schemas containing millions of entities without significant performance degradation [84].
@@
On the one hand, LLMs have attracted significant attention from researchers in exploring their potentials for various scenarios of IE. In addition to excelling in individual IE tasks, LLMs possess a remarkable ability to effectively model various IE tasks in a universal format. This is conducted by capturing inter-task dependencies with instructive prompts, and achieves consistent performance [133, 175, 160, 204, 52, 47, 111]. On the other hand, recent works have shown the outstanding generalization of LLMs to not only learn from IE training data through fine-tuning [160, 223, 79, 16, 160], but also extract information in few-shot and even zero-shot scenarios relying solely on in-context examples or instructions [210, 205, 203, 5, 220]. However, for above two groups of research works: 1) the universal frameworks for multiple tasks; 2) cutting-edge learning techniques for scenarios with limited training data, existing surveys [149, 247, 225] do not provide a comprehensive exploration of these areas.
In this survey, we provide a comprehensive exploration of LLMs for generative IE, as illustrated in Figure 1. To achieve this, we categorize existing methods mainly using two taxonomies: (1) a taxonomy of numerous IE subtasks, which aims to classify different types of information that can be extracted individually or uniformly, and (2) a taxonomy of IE techniques, which categorizes various novel approaches that utilize LLMs for generative IE, particularly on low-resource scenarios. In addition, we present a comprehensive review of studies that specifically focus on the application of IE techniques in various domains. And we discuss studies that aim to evaluate and analyze the performance of LLMs for IE. We also compare several representative methods to gain deeper understanding of their potentials and limitations, and provide insightful analysis on future directions. To the best of our knowledge, this is the first survey on generative IE with LLMs.
@@
& Typing, leaf-task, text width=5.5em [Yan et al. [223], TEMPGEN [79], Cui et al. [31], Zhang et al. [237], Wang et al. [201], Xia et al. [215], Cai et al. [17],
EnTDA [73], Amalvy et al. [4], GPT-NER [203], Cp-NER [26], LLMaAA [236], PromptNER [5], Ma et al. [140],
Xie et al. [218], UniNER [249], NAG-NER [239],  Su and Yu [182], GNER [36], NuNER [13], MetaNER [22],
LinkNER [241], SLCoLM [188], C-ICL [143],  Popovič and Färber [162],  Keloth et al. [88], ProgGen [69], VerifiNER [92],
 Li et al. [115],  Oliveira et al. [154],  Lu et al. [131],  Bölücü et al. [14],  Liu et al. [121], ConsistNER [213],  Naguib et al. [147],
GLiNER [230],  Munnangi et al. [146],  Zhang et al. [234], LTNER [222], ToNER [82],  Nunes et al. [153], etc. , modelnode-task, text width=45em] ] ] [ Relation
@@
 Otto et al. [156],  Shi and Luo [179],  Li et al. [102],  Li et al. [104], RAG4RE [44],  Li et al. [114], modelnode-task, text width=45em] ] [ Triplet, leaf-task, text width=5.5em [TEMPGEN [79], Fan and He [46], Kwak et al. [95], DocRTE [183], REPLM [158], AutoRE [117], ERA-COT [126],
@@
Extraction, leaf-task, text width=5.5em [BART-Gen [112], Text2Event [132], ClarET [250], X-GEAR [77], PAIE [139], GTEE-DYNPREF [125], Cai and O’Connor [18],
@@
Extraction, leaf-task, text width=5.5em [BART-Gen [112], Text2Event [132], Hsu et al. [71], ClarET [250], GTEE-DYNPREF [125], Ma et al. [140], GL [159],
STAR [137], Cai and O’Connor [18], DemoSG [242],  Kwak et al. [95], EventRL [54],  Huang et al. [78], modelnode-task,text width=45em] ] ] [ Universal
@@
ntation (§4.1), leaf-paradigms, text width=5.5em [ Data Annotation, leaf-paradigms,text width=9em [Veyseh et al. [193], LLMaAA [236], AugURE [202], Xu et al. [220], Li et al. [109], Tang et al. [186], REPAL [248],
Meoni et al. [142],  Evans et al. [45], MetaIE [161], NuNER [13], LLM-DA [226],  [185],  [148],  [154], modelnode-paradigms,text width=41.5em] ] [ Knowledge Retrieval, leaf-paradigms, text width=9em [Amalvy et al. [4], Chen and Feng [21], PGIM [107], DocRTE [183], VerifiNER [92], RiVEG [108], modelnode-paradigms,text width=41.5em] ] [ Inverse Generation, leaf-paradigms, text width=9em [EnTDA [73], STAR [137], QGA-EE [130], SynthIE [85], ProgGen [69],  Atuhurra et al. [6], modelnode-paradigms,text width=41.5em] ] [ Synthetic Datasets for
Instruction-tuning , leaf-paradigms, text width=9em [UniversalNER [249], GLiNER [230], GNER [36], Chen and Feng [21], modelnode-paradigms,text width=41.5em] ] ] [ Prompts Design
4.2), leaf-paradigms, text width=5.5em [ Question Answer, leaf-paradigms, text width=9em [QA4RE [233], ChatIE [210],  Hou et al. [70],  Otto et al. [156],  Li et al. [100],modelnode-paradigms, text width=41.5em] ] [ Chain of Thought, leaf-paradigms,text width=9em [ Chen and Feng [21], PromptNER [5],  Wadhwa et al. [194],  Yuan et al. [228],  Bian et al. [10], RT [110], modelnode-paradigms,text width=41.5em] ] [ Self-Improvement, leaf-paradigms, text width=9em [ Xie et al. [218], ProgGen [69], ULTRA [238],modelnode-paradigms, text width=41.5em] ] ] [ Zero-shot
@@
Prompting, leaf-paradigms, text width=6em [Xie et al. [218], QA4RE [233], Cai and O’Connor [18], AugURE [202], Li et al. [100], Code4Struct [205], CodeKGC [9],
@@
 Li et al. [103],  Li et al. [104],  Li et al. [114],  Sun et al. [185],  Hu et al. [75],  Shao [178], modelnode-paradigms,text width=44.5em] ] [ Cross-Domain
@@
TEXTEE [78],  Kwak et al. [95],  Bölücü et al. [14],  Qi et al. [165],  Zhang et al. [234],  Naguib et al. [147],
@@
 Su and Yu [182],  Zaratiana et al. [231],  Fan and He [46],  Popovič and Färber [162],  Keloth et al. [88], GLiNER [230],  Ding et al. [37], etc. , modelnode-paradigms,text width=52em] ] ] [ Specific
@@
5), leaf-others,text width=5em [ Multimodal (Chen and Feng [21] PGIM [107] Cai et al. [17] RiVEG [108]), Multilingual (X-GEAR [77] Meoni et al. [142] Naguib et al. [147]), Scientific (
Dunn et al. [42] PolyIE [27] Bölücü et al. [14] Foppiano et al. [50] Dagdelen et al. [32]), Medical (DICE [136] Tang et al. [186] Hu et al. [74] Meoni et al. [142]
Bian et al. [10] RT [110] Agrawal et al. [2] Labrak et al. [96] Gutiérrez et al. [66] VANER [12] Munnangi et al. [146] Monajatipoor et al. [144] Hu et al. [75]),
Astronomical (Shao [178] Evans et al. [45]), Historical (González-Gallardo et al. [58] CHisIEC[187]), Legal (Nunes et al. [153] Oliveira et al. [154] Kwak et al. [95]), modelnode-others, text width=59em] ] [ Evaluation
@@
6), leaf-others,text width=5em [Gutiérrez et al. [66], GPT-3+R [2], Labrak et al. [96], Xie et al. [217], Gao et al. [53], Li et al. [110], InstructIE [60], Han et al. [67], Katz et al. [87], Hu et al. [74],
González-Gallardo et al. [58], Yuan et al. [228], Wadhwa et al. [194], PolyIE [27], Li et al. [98], Qi et al. [166], XNLP [48], IEPile [61], Naguib et al. [147],
@@ 2 Preliminaries of Generative IE
In this section, we provide a formal definition of discriminative and generative IE and summarize the IE subtasks, as outlined in Ye et al. [225]. This survey focuses primarily on the tasks of Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE) [204, 175], as these are tasks that receive the most attention in IE papers.
@@ 3 LLMs for Different Information Extraction Tasks
3.1 Named Entity Recognition
NER is a crucial component of IE and can be seen as a predecessor or subtask of RE and EE. It is also a fundamental task in other NLP tasks, thus attracting significant attention from researchers to explore new possibilities in the era of LLMs [70, 110, 81, 45, 11, 144, 156, 148, 161, 6, 187, 226, 108, 75, 151, 152, 50, 32, 178]. Considering the gap between the sequence labeling and generation models, GPT-NER [203] transformed NER into a generative task and proposed a self-verification strategy to rectify the mislabeling of NULL inputs as entities.
@@ 3.1 Named Entity Recognition
Table 1: Comparison of LLMs for named entity recognition (identification & typing) with the Micro-F1 metric (%). indicates that the model is discriminative. We demonstrate some universal and discriminative models for comparison. IE techniques include Cross-Domain Learning (CDL), Zero-Shot Prompting (ZS Pr), In-Context Learning (ICL), Supervised Fine-Tuning (SFT), Data Augmentation (DA). Uni. denotes whether the model is universal. Onto. 5 denotes the OntoNotes 5.0. Details of datasets (§8) and backbones (§8) are presented in Appendix. The settings for all subsequent tables are consistent with this format.
Xie et al. [218] proposed a training-free self-improving framework that uses LLM to predict on the unlabeled corpus to obtain pseudo demonstrations, thereby enhancing the performance of LLM on zero-shot NER.
@@
Table 1 shows the comparison of NER on five main datasets, which are obtained from their original papers. We can observe that: 1) the models in few-shot and zero-shot settings still have a huge performance gap behind the SFT and DA. 2) Even though there is little difference between backbones, the performance varies greatly between methods under the ICL paradigm. For example, GPT-NER opens up at least a 6% F1 value gap with other methods on each dataset, and up to about 19% higher. 3) Compared to ICL, there are only minor differences in performance between different models after SFT, even though the parameters in their backbones can differ by up to a few hundred times.
3.2 Relation Extraction
RE also plays an important role in IE, which usually has different setups in different studies as mentioned in Section 2. To address the poor performance of LLMs on RE tasks due to the low incidence of RE in instruction-tuning datasets, as indicated by Gutiérrez et al. [66], QA4RE [233] introduced a framework that enhances LLMs’ performance by aligning RE tasks with QA tasks. GPT-RE [196] incorporates task-aware representations and enriching demonstrations with reasoning logic to improve the low relevance between entity and relation and the inability to explain input-label mappings. Due to the large number of predefined relation types and uncontrolled LLMs, Li et al. [109] proposed to integrate LLM with a natural language inference module to generate relation triples, enhancing document-level relation datasets.
As shown in the Table 2 and 3, we statistically found that universal IE models are generally better solving harder Relation Strict problems due to learning the dependencies between multi-tasks [160, 133], while the task-specific methods solve simpler RE subtasks (e.g. relation classification). In addition, compared with NER, it can be found that the performance differences between models in RE are more obvious, indicating that the potential of LLM in RE task.
3.3 Event Extraction
Events can be defined as specific occurrences or incidents that happen in a given context. Recently, many studies [125, 130] aim to understand events and capture their correlations by extracting event triggers and arguments using LLMs, which is essential for various reasoning tasks [8]. For example, Code4Struct [205] leveraged LLMs to translate text into code to tackle structured prediction tasks, using programming language features to introduce external knowledge and constraints through alignment between structure and code. Considering the interrelation between different arguments in the extended context, PGAD [135] employed a text diffusion model to create a variety of context-aware prompt representations, enhancing both sentence-level and document-level event argument extraction by identifying multiple role-specific argument span queries and coordinating them with the context.
3.4 Universal Information Extraction
Code-LLMs. On the other hand, code-based methods unify IE tasks by generating code with a universal programming schema [205]. Code4UIE [64] proposed a universal retrieval-augmented code generation framework, which leverages Python classes to define schemas and uses in-context learning to generate codes that extract structural knowledge from texts. Besides, CodeKGC [9] leveraged the structural knowledge inherent in code and employed schema-aware prompts and rationale-enhanced generation to improve performance. To enable LLMs to adhere to guidelines out-of-the-box, GoLLIE [175] enhanced zero-shot ability on unseen IE tasks by aligning with annotation guidelines.

### v3

@@ Abstract
[…]
@@
†Equal Contribution. The article has been accepted by Frontiers of Computer Science (FCS), with the DOI: 10.1007/s11704-024-40555-y. You can cite the FCS version.,
@@ 1 Introduction
Information Extraction (IE) is a crucial domain in natural language processing (NLP) that converts plain text into structured knowledge (e.g., entities, relations, and events), and serves as a foundational requirement for a wide range of downstream tasks, such as knowledge graph construction [1], knowledge reasoning [2] and question answering [3]. Typical IE tasks consist of Named Entity Recognition (NER), Relation Extraction (RE) and Event Extraction (EE) [4, 5, 6, 7]. However, performing IE tasks is inherently challenging. This is because IE involves extracting information from various sources and dealing with complex and ever-changing domain requirements [8]. Unlike traditional NLP tasks, IE encompasses a broad spectrum of objectives such as entity extraction, relationship extraction, and more. In IE, the extraction targets exhibit intricate structures where entities are presented as span structures (string structures) and relationships are represented as triple structures [4]. Additionally, in order to effectively handle different information extraction tasks, it is necessary to employ multiple independent models. These models are trained separately for each specific task, without sharing any resources. However, this approach comes with a drawback: managing a large number of information extraction models becomes costly in terms of the resources needed for construction and training, like annotated corpora.
The emergence of large language models (LLMs), such as GPT-4 [9], has significantly advanced the field of NLP, due to their extraordinary capabilities in text understanding and generation [10, 11, 12]. Pretraining LLMs using auto-regressive prediction allows them to capture the inherent patterns and semantic knowledge within text corpora [13, 14, 15, 16, 17, 18, 19]. This enhances LLMs with the capability to perform zero-shot and few-shot learning, enabling them to model various tasks consistently and serving as tools for data augmentation [20, 21, 22]. Furthermore, LLMs can serve as intelligent agents for complex task planning and execution, utilizing memory retrieval and various tools to enhance efficiency and successfully accomplish tasks [23, 24, 25, 26, 27]. Therefore, there has been a recent surge of interest in generative IE methods [28] that adopt LLMs to generate structural information rather than extracting structural information from plain text. These methods have been proven to be more practical in real-world scenarios compared to discriminated methods [29, 30], as they can handle schemas containing millions of entities without significant performance degradation [31].
@@
On the one hand, LLMs have attracted significant attention from researchers in exploring their potentials for various scenarios and tasks of IE. In addition to excelling in individual IE tasks, LLMs possess a remarkable ability to effectively model all IE tasks in a universal format. This is conducted by capturing inter-task dependencies with instructive prompts, and achieves consistent performance [4, 32, 33, 5, 34, 35, 36]. On the other hand, recent works have shown the outstanding generalization of LLMs to not only learn from IE training data through fine-tuning [33, 37, 38, 39, 33], but also extract information in few-shot and even zero-shot scenarios relying solely on in-context examples or instructions [40, 41, 42, 43, 44].
However, existing surveys [45, 8, 46] do not provide a comprehensive exploration of these areas for the above two groups of research works: 1) universal frameworks that cater to multiple tasks and 2) cutting-edge learning techniques for scenarios with limited training data. The community urgently needs a more in-depth analysis of how LLM can be more appropriately applied to IE tasks to improve the performance of the IE field. This is because there are still challenges and issues in applying LLM to IE in terms of learning and understanding knowledge [47]. These challenges include the misalignment between natural language output and structured form [6], hallucination problem in LLMs [48], contextual dependence, high computational resource requirements [49], difficulties in updating internal knowledge [50], etc.
In this survey, we provide a comprehensive exploration of LLMs for generative IE, as illustrated in Fig. 1. To achieve this, we categorize existing methods mainly using two taxonomies: (1) a taxonomy of numerous IE subtasks, which aims to classify different types of information that can be extracted individually or uniformly, and (2) a taxonomy of IE techniques, which categorizes various novel approaches that utilize LLMs for generative IE, particularly on low-resource scenarios. In addition, we present a comprehensive review of studies that specifically focus on the application of IE techniques in various domains. And we discuss studies that aim to evaluate and analyze the performance of LLMs for IE. According to the above division, we construct a taxonomy of related studies as shown in Fig. 2. We also compare several representative methods to gain deeper understanding of their potentials and limitations, and provide insightful analysis on future directions. To the best of our knowledge, this is the first survey on generative IE with LLMs.
@@
& Typing, leaf-task, text width=5.5em [Yan et al. [37], TEMPGEN [38], Cui et al. [53], Zhang et al. [54], Wang et al. [55], Xia et al. [56], Cai et al. [57], EnTDA [58],
Amalvy et al. [59], GPT-NER [42], Cp-NER [60], LLMaAA [61], PromptNER [43], Ma et al. [62], Xie et al. [63], UniNER [64],
NAG-NER [65], Su et al. [66], GNER [67], NuNER [68], MetaNER [69], LinkNER [70], SLCoLM [71], Popovivc et al. [72],
ProgGen [73], C-ICL [74], Keloth et al. [75], VerifiNER [76], Li et al. [77], Oliveira et al. [78], Lu et al. [79], Bolucu et al. [80],
Liu et al. [81], ConsistNER [82], Naguib et al. [83], GLiNER [84], Munnang et al. [85], Zhang et al. [86], LTNER [87],
ToNER [88], Nunes et al. [89], Hou et al. [90], RT [91], Jiang et al. [92], VANER [93], RiVEG [94], LLM-DA [95], etc. , modelnode-task, text width=45em] ] ] [ Relation
@@
Otto et al. [108], Shi et al. [109], Li et al. [110], Li et al. [111], RAG4RE [112], Li et al. [113], modelnode-task, text width=45em] ] [ Triplet, leaf-task, text width=5.5em [TEMPGEN [38], Fan et al. [114], Kwak et al. [115], DocRTE [116], REPLM [117], AutoRE [118], ERA-COT [119],
@@
Extraction, leaf-task, text width=5.5em [BART-Gen [130], Text2Event [131], ClarET [132], X-GEAR [133], PAIE [134], GTEE-DYNPREF [135], Cai et al. [136],
@@
Extraction, leaf-task, text width=5.5em [BART-Gen [130], Text2Event [131], DEGREE [147], ClarET [132], GTEE-DYNPREF [135], Ma et al. [62], GL [97],
STAR [100], Cai et al. [136], DemoSG [148], Kwak et al. [115], EventRL [149], Huang et al. [150], modelnode-task,text width=45em] ] ] [ Universal
@@
ntation (§4.1), leaf-paradigms, text width=5.5em [ Data Annotation, leaf-paradigms,text width=9em [Veyseh et al. [127], LLMaAA [61], AugURE [101], Xu et al. [44], Li et al. [161], Tang et al. [162], REPAL [105],  [78]
Meoni et al. [163], Evans et al. [164], MetaIE [124], NuNER [68], LLM-DA [95], Sun et al. [145], Naraki et al. [165], modelnode-paradigms,text width=41.5em] ] [ Knowledge Retrieval, leaf-paradigms, text width=9em [Amalvy et al. [59], [166], PGIM [167], DocRTE [116], VerifiNER [76], RiVEG [94], modelnode-paradigms,text width=41.5em] ] [ Inverse Generation, leaf-paradigms, text width=9em [EnTDA [58], STAR [100], QGA-EE [138], SynthIE [168], ProgGen [73], Atuhurra et al. [125], modelnode-paradigms,text width=41.5em] ] [ Synthetic Datasets for
Instruction-tuning , leaf-paradigms, text width=9em [UniversalNER [64], GLiNER [84], GNER [67], Chen et al. [166], modelnode-paradigms,text width=41.5em] ] ] [ Prompts Design
4.2), leaf-paradigms, text width=5.5em [ Question Answer, leaf-paradigms, text width=9em [QA4RE [98], ChatIE [40], Hou et al. [90], Otto et al. [108], Li et al. [96],modelnode-paradigms, text width=41.5em] ] [ Chain of Thought, leaf-paradigms,text width=9em [Chen et al. [166], PromptNER [43], Wadhwa et al. [169], Yuan et al. [170], Bian et al. [171], RT [91], modelnode-paradigms,text width=41.5em] ] [ Self-Improvement, leaf-paradigms, text width=9em [Xie et al.  [63], ProgGen [73], ULTRA [144],modelnode-paradigms, text width=41.5em] ] ] [ Zero-shot
@@
Prompting, leaf-paradigms, text width=6em [Xie et al. [63], QA4RE [98], Cai et al. [136], AugURE [101], Li et al. [96], Code4Struct [41], CodeKGC [159],
@@
Li et al. [107], Li et al. [111], Li et al. [113], Sun et al. [145], Hu et al. [172], Shao et al. [173], modelnode-paradigms,text width=44.5em] ] [ Cross-Domain
@@
TEXTEE [150], Kwak et al. [115], Bolucu et al. [80], Qi et al. [106], Zhang et al. [86], Naguib et al. [83],
@@
Su et al. [66], Zaratiana et al. [123], Fan et al. [114], Popovivc et al. [72], Keloth et al. [75], GLiNER [84], Ding et al. [121], etc. , modelnode-paradigms,text width=52em] ] ] [ Specific
@@
5), leaf-others,text width=5em [ Multimodal (Chen et al. [166] PGIM [167] Cai et al. [57] RiVEG [94]), Multilingual (X-GEAR [133] Meoni et al. [163] Naguib et al. [83]), Scientific (Dunn et al. [180]
PolyIE [181] Bolucu et al. [80] Foppiano et al. [47] Dagdelen et al. [182]), Medical (DICE [183] Tang et al. [162] Hu et al. [184] Meoni et al. [163] Bian et al. [171]
RT [91] Agrawal et al. [185] Labrak et al. [186] Gutierrez et al. [187] VANER [188] Munnangi et al. [85] Monajatipoor et al. [179] Hu et al. [172]),
Astronomical (Shao et al. [173] Evans et al. [164]), Historical (Gonzalez et al. [189] CHisIEC[126]), Legal (Nune et al. [89] Oliveira et al. [78] Kwak et al. [115]), modelnode-others, text width=59em] ] [ Evaluation
@@
6), leaf-others,text width=5em [Gutierrez et al. [187], GPT-3+R [185], Labrak et al. [186], Xie et al. [190], Gao et al. [191], RT [91], InstructIE [192], Han et al. [193], Katz et al. [194], Hu et al. [184],
Gonzalez et al. [189], Yuan et al. [170], Wadhwa et al. [169], PolyIE [181], Li et al. [195], Qi et al. [28], XNLP [196], IEPile [157], Naguib et al. [83],
@@ 2 Preliminaries of Generative IE
In this section, we provide a formal definition of discriminative and generative IE and summarize the IE subtasks, as outlined in [46]. This survey focuses primarily on the tasks of Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE) [5, 32], as these are tasks that receive the most attention in IE papers. Examples are shown in Fig. 3.
@@ 3 LLMs for Different Information Extraction Tasks
Table 1: Comparison of LLMs for named entity recognition (identification and typing) with the Micro-F1 metric (%). indicates that the model is discriminative. We demonstrate some universal and discriminative models for comparison. IE techniques include Cross-Domain Learning (CDL), Zero-Shot Prompting (ZS Pr), In-Context Learning (ICL), Supervised Fine-Tuning (SFT), Data Augmentation (DA). Uni. denotes whether the model is universal. Onto. 5 denotes the OntoNotes 5.0. Details of datasets and backbones are presented in Section 8. The settings for all subsequent tables are consistent with this format.
@@
YAYI-UIE [155] SFT $\surd$ Baichuan2-13B - 81.78 96.77 87.04 75.21
ToNER [88] SFT Flan-T5-3B 88.09 86.68 93.59 91.30 -
KnowCoder [160] SFT $\surd$ LLaMA2-7B 86.2 86.1 95.1 88.2 76.7
GNER [67] SFT Flan-T5-11B - - 93.28 91.83 -
@@
3.1 Named Entity Recognition
NER is a crucial component of IE and can be seen as a predecessor or subtask of RE and EE. It is also a fundamental task in other NLP tasks, thus attracting significant attention from researchers to explore new possibilities in the era of LLMs [90, 91, 92, 164, 93, 179, 108, 165, 124, 125, 126, 95, 94, 172, 122, 177, 47, 182, 173]. Considering the gap between the sequence labeling and generation models, GPT-NER [42] transformed NER into a generative task and proposed a self-verification strategy to rectify the mislabeling of NULL inputs as entities. Xie et al. [63] proposed a training-free self-improving framework that uses LLM to predict on the unlabeled corpus to obtain pseudo demonstrations, thereby enhancing the performance of LLM on zero-shot NER.
Table 1 shows the comparison of NER on five main datasets, which are obtained from their original papers. We can observe that:
  • 1) the models in few-shot and zero-shot settings still have a huge performance gap behind the SFT and DA.
@@ 3.1 Named Entity Recognition
  • 2) Even though there is little difference between backbones, the performance varies greatly between methods under the ICL paradigm. For example, GPT-NER opens up at least a 6% F1 value gap with other methods on each dataset, and up to about 19% higher.
  • 3) Compared to ICL, there are only minor differences in performance between different models after SFT, even though the parameters in their backbones can differ by up to a few hundred times.
  • 4) The performance of models trained with the SFT paradigm exhibits greater variability across datasets, particularly for universal models. For instance, YAYI-UIE [155] and KnowCoder [160] outperform other models by at least 2.89% and 1.22% respectively on CoNLL03, while experiencing a decrease of 7.04% and 5.55% respectively compared to the best model on GENIA. We hypothesize that this discrepancy may arise from these models being trained on diverse datasets primarily sourced from news and social media domains, whereas GENIA represents a smaller fraction in the training set as it belongs to the biomedical domain; thus resulting in significant distribution gaps between different fields that ultimately impact performance outcomes. Furthermore, universal models necessitate simultaneous training across multiple subtasks, which inevitably exacerbates this distribution gap.
  • 5) The EnTDA [58], on the contrary, exhibits exceptional stability and outperforms other methods on all datasets, thereby substantiating the robustness of the DA paradigm in addressing specific tasks.
@@
YAYI-UIE[155] SFT $\surd$ Baichuan2-13B 89.97 - 84.41 79.73 40.94
KnowCoder [160] SFT $\surd$ LLaMA2-7B 93.7 64.5 84.8 73.3 40.0
@@
3.2 Relation Extraction
RE also plays an important role in IE, which usually has different setups in different studies as mentioned in Section 2. To address the poor performance of LLMs on RE tasks due to the low incidence of RE in instruction-tuning datasets, as indicated by [187], QA4RE [98] introduced a framework that enhances LLMs’ performance by aligning RE tasks with QA tasks. GPT-RE [99] incorporates task-aware representations and enriching demonstrations with reasoning logic to improve the low relevance between entity and relation and the inability to explain input-label mappings. Due to the large number of predefined relation types and uncontrolled LLMs, Li et al. [161] proposed to integrate LLM with a natural language inference module to generate relation triples, enhancing document-level relation datasets.
As shown in the Table 2 and 3, we statistically found that universal IE models are generally better solving harder Relation Strict problems due to learning the dependencies between multi-tasks [33, 4], while the task-specific methods solve simpler RE subtasks (e.g. relation classification). Moreover, when compared to NER, it becomes evident that the performance disparities among models in RE are more pronounced, thereby highlighting the potential of LLM in addressing RE tasks.